# Photoluminescence from metal nanostructures - dependence on size

### v1

@@ 1 Introduction
The emission of light from metal nanostructures induced by illumination, frequently referred to colloquially as metal photoluminescence (PL), is a fundamental process occurring in a prototypical solid-state system111In many cases, the term photoluminescence is used to refer to spontaneous emission / radiative recombination of an electron and a hole, in contrast to inelastic light scattering, aka electronic Raman scattering. In that sense, in this work we use the term photoluminescence (PL) as a shorthand for emission, without distinguishing between these possibilities.. It also has practical importance - metal nanoparticles are frequently used as non-bleaching fluorescent labels in bio-imaging 1, 2, correlative light-electron microscopy 3, and more recently, as enablers of thermometry 4, 5, 6, 7, 8.
@@
Despite that dual importance, it is surprising that many aspects of the metal PL have been under debate for decades. These debates are on, for example, whether the emission occurs due to recombination of an electron and hole residing both within the conduction band 9, 10, 11 or involving a hole in the valence band (e.g., Ref. [ 12, 13, 14, 15, 16, 17]), on whether the emission statistics is fermionic or bosonic (e.g., Ref. [ 6, 18]), thermal or non-thermal (e.g., Ref. [ 19, 20, 10, 21, 11]), or how it depends on the local electric field (e.g., Ref. [ 19, 20, 10, 22]). Of particular interest is the dependence on the size of the nanostructure (e.g., Ref. [ 23, 24, 25, 15, 16]). In particular, many early works studied the emission from metal clusters or few nm particles, primarily focussing on atomistic considerations and ligand coverage; these works highlighted the trade-off between the presence of a plasmonic resonance and the quantum yield for emission, see Ref. [ 26]. For larger structures, there have been reports of contradicting trends. For example, Dulkeith et al. Ref. [ 23] showed that the (time and) frequency integrated Stokes emission (SE) from spheres of growing radii illuminated by a short pulse scales linearly with their volume; similar results were reported by Gaiduk et al. Ref. [ 24] for continuous wave (CW) illumination. On the other hand, Lin et al. Ref. [ 27] and more recently, Bowman et al. Ref. [ 28] have demonstrated that the CW PL from rods and thin Au films, respectively, decreases with growing thickness; similar trends were reported in Ref. [ 16, 29] for pulsed illumination. Other works reported a varying trend (e.g., Ref. [ 25]).
@@
These debates originate from the different conditions under which the PL measurements have been made (in terms of the number, density and shape of the nanostructure geometry, the illumination duration, intensity and wavelength, the relative position of the latter with respect to the resonances of the structure, the relative spectral position of emission with respect to the illumination wavelength, the geometry and thermal properties of the surrounding etc.) or the exact quantity studied, but also from the absence of a simple theory that encompasses both photonic aspects of the problem (via the photonic density of states, as appearing in Purcell’s formula), as well as the thermal aspects and the electronic aspects, i.e., the distribution and associated weights of the continuum of possible recombination transitions.
@@
Recently, Dubi and Sivan employed a simple analytic expression for the steady-state electron distribution in a Drude metal to derive an equally simple expression for the PL from metals illuminated by CW light 11. By relying on the empiric values of the permittivity to determine the absorption, this approach circumvents the need to specify the transition matrix element and circumvents the arguments about the origins of the emission (being radiative recombination or electronic Raman / inelastic light scattering etc. 4, 30, 28, 31, 32, 33). This expression enabled the resolution of many of the disagreements described above (in particular, associated with the emission statistics and electric-field dependence). This expression was more recently extended to explain the dependence of the PL following short pulse illumination on the electric field 34.
@@
In this work, we use the approach of Ref. [ 11] to reconcile the various seemingly contradicting reports on the size-dependence of the PL from the prototypical geometries of metal particles and layers. Specifically, in Section 2, we combine the analytic expression of Ref. [ 11] with the detailed analysis of the heating of illuminated metal spheres 35 and most importantly, with recent progress made by Loirette-Pelous and Greffet who showed how to determine the total PL of a metal nanostructure in the case of a non-thermal distribution 36.
@@
In Section 3, we use the resulting analytic expression to compute the PL from metal nano-spheres and nano-layers as a function of their size. We distinguish between three cases. For weak illumination (hence, negligible heating), the dependence of the PL spectrum (both the Stokes Emission (SE) and the anti-Stokes Emission (aSE)) on the structure size is determined by a single parameter - the absorption cross-section density (or equivalently, the emission cross-section density). Thus, generically, the PL scales with the illuminated volume - it is a volume effect for small sizes and it becomes a surface effect when the size exceeds the light penetration depth. In the latter case, the resonances characteristic of particles modify further the size-dependence of the PL. Indeed, a change of size causes the emission at a given frequency to shift in and out of resonance.
@@
For stronger illumination (hence, moderate heating), the size-dependence of the PL is amplified through the dependence of the Bose function on the (electron) temperature. This has a fairly small effect on the SE, but a large effect on the aSE due to the exponential dependence of the Bose function on the temperature, which itself scales with the absorption cross-section density. For even stronger heating, thermo-optic effects kick in, and cause the quality factor of the resonance to decrease 37, 38. This causes weaker (excessive) heating at resonance (away from resonance), and hence, has a complex effect on the PL, depending on the emission frequency (SE/aSE).
@@
We then demonstrate a good qualitative match between the prediction of our model and experimental results for spheres and layers, and even good quantitative matches in most cases. We demonstrate different trends by looking also at other particles shapes (rods and nano disks), and find a good qualitative match, even in the presence of modest field and temperature non-uniformity levels. Remarkably, the agreement extends beyond the formal limits of the analysis, specifically, in the presence of interband emission events, which are not accounted for in our analytic expression.
@@ 2.1 A microscopic view
In Ref. [ 11], Sivan and Dubi presented a quantitative theory for the PL from Drude metals under continuous wave (CW) illumination, showing that the probability of emitted photons per unit frequency is given approximately by
@@
represents the emission rate of a single electron. In the above, and are the emitter position vector and frequency, is the excitation frequency, is the Fermi energy, is the electron density of states and is the local field. is the non-equilibrium electron distribution function, is the transition dipole moment between electronic states with an initial energy and final energy (assumed to be energy-indepdendent) and is the local density of photonic states (LDOPS).
Based on the CW solution for (the steady-state electron distribution) obtained in Ref. [ 39, 40] for uniformly illuminated Drude metals and its experimental verification in Ref. [ 41, 42, 43, 44], the electronic part of the emission formula in Eq. (2) was shown in Ref. [ 11] to consist of a series of Planck’s black-body radiation-like terms , i.e.,
@@
Here, is the electron temperature; for CW illumination, the latter is nearly equal to the phonon temperature, however, since the PL literature is inconsistent on the matter, the notation adopted in this manuscript emphasizes that the dependence is on the electron temperature. Further, the absorbed power density () is defined in terms of the local as
@@
where is the electron density and is an average rate of collisions between electrons. For simplicity, we set in Eq. (7) to a value typical for a non-thermal electron, i.e., at .
The first term in Eq. (2.1), represents the average energy of thermal emission per electromagnetic mode (i.e., for vacuum electric fields); the next terms represent the non-thermal emission induced by deviations of the electron distribution from thermal equilibrium due to one photon absorption (1PA), two photon absorption (2PA) etc.. As discussed in Ref. [ 11, 34], for CW illumination, typically these terms are small compared to the thermal emission at mid-IR frequencies, but dominate the emission close to the illumination frequencies and higher 11, 34 as depicted in the schematic representation (Fig. 1). The complete step structure of the non-thermal contributions was observed experimentally for CW illumination for the first time in Ref. [ 17].
Notably, the emission (1) is assumed to occur strictly within the conduction band, i.e., to involve only intraband transitions. In the presence of interband transitions, this expression can be amended by accounting for the increased number of non-thermal holes above the interband absorption threshold, as shown in Fig. S10(b) of Ref. [ 28]; as shown below, our analysis provides a qualitative (and usually, also a quantitative) match to the experimental data even without accounting for these additional transitions.
@@ 2.2 A macroscopic view
The result (1) applies for a general point within the metal. It was so far used to understand the general parametric dependence of the PL on the electric field, its statistics etc. with only a phenomenological address of the shape and size of the nanostructure.
@@
In order to account for a specific nanostructure geometry, one needs to sum over the random spontaneous emission events from the nanostructure volume as a function of space. This should allow one to account also for the actual magnitude of the emission, thus, including also the portion of the emitted photons that got re-absorbed (sometimes referred to as re-cycled), rather than only the probability of emission (as given by (1)). This can be done using the local Kirchhoff Law for non-isothermal bodies 45, 46, 47 and its extension to metals having non-equilibrium electron distributions 36. The latter study showed that the actual emitted power in the direction into a solid angle (; see Eq. (9.15.9) in Ref. [ 48]) is given by
where is the direction of incidence, and are the polarization of the emitted and incident waves, respectively, and is the emission cross-section density (aka emissivity density), which by the local Kirchhoff Law, equals the absorption cross-section density. Unlike in Ref. [ 36], where the PL was expressed in terms of absorption cross-section, Eq. (8), utilizes the local absorption cross-section density to extend the formulation to structures with nonuniform field distributions, similar to the approach used in Ref. [ 45, 46, 47]. It is notable to mention again that since the formula above accounts only for emission within the conduction band, then, strictly, the emissivity density represented by should include only the intraband contribution to the absorption cross-section density (as in Ref. [ 36]). This would not matter for the illustrative examples below, for which we study Ag nanostructures, but could somewhat affect the match between our predictions and the experimental data (see Sections 3.1.5 and 3.2.3).
@@
Eq. (9) shows that the quantities that determine the emission are the absorption cross-section density and the (electron) temperature distribution. As shown below, the simplified expression (9) also circumvents the more advanced yet complicated -space calculations (as e.g., in Ref. [ 29, 28, 32]), without compromising much predictive capabilities.
@@
We now note that the spatially-integrated emission (9) involves space-varying functions. Specifically, the spatial distribution of depends on the material constituents and geometry of the nanostructure. In addition to gradients induced by illumination non-uniformity (e.g., when a focussed beam illuminates a film), it exhibits significant gradients on scales exceeding the penetration (skin) depth, i.e., for more than a few tens of nm for noble metals. In contrast, the non-uniformity of temperature is typically much weaker than that of the electric field (and hence ) due to the strong (electron, hence) heat diffusion in metals. Specifically, the (electron) temperature non-uniformity should, in principle, be extracted from a self-consistent solution of microscopic equations for the electron dynamics; however, in practice, since the deviation from thermal equilibrium is minute for CW illumination 39, 40, 44, it is usually determined by coarse-graining such equations into heat equations, see, e.g., Ref. [ 49, 50, 39, 40]. Accordingly, the derived single or two temperature models reveal that heat gradients and electron-phonon temperature differences are very small (see, e.g., Ref. [ 51, 52, 53, 54, 55, 56]). Finally, like the absorption cross-section density and unlike the electron temperature , the non-thermal part of the electron distribution, manifested via (5), is roughly determined by the local electric field distribution. As predicted in Ref. [ 30, 57, 58] and demonstrated experimentally in Ref. [ 28], the reason for that is the minimal (few nm) mean free path of these electrons, which in turn, originates from their femtosecond-scale collision rate. Accordingly, in what follows we will use heat equations to determine the electron temperature and rely on the solution of Maxwell’s equations to determine the electric field, and hence, the spatial distribution of and . Having said that, in many cases (specifically, for few nm spheres and uniformly illuminated thin films, both studied below), is quite uniform. In such cases, can be estimated from measurable quantities like the absorption cross-section, (in the case of particles) or absorptance, (in the case of films), thus, simplifying the formulation slightly, and making the analysis of the PL simpler.
@@ 3.1.1 Determination of the electron temperature
We start our analysis by considering the PL from nano-spheres under CW plane wave illumination in a uniform optical and thermal environment (oil, in our case), see Fig. 2(a). We choose silver as a prototypical plasmonic material, as it does not require accounting for interband transitions for illumination with visible light. As mentioned above, in this case, it is customary to neglect the small difference between the electron and phonon temperatures (see, e.g., Ref. [ 59, 39, 40]) and consider a single temperature model. Nevertheless, as mentioned, we choose to denote the temperature below as , in order to emphasize that it is the electron temperature that is the relevant quantity as far as photon emission is concerned. Then, the steady-state temperature can be obtained by solving the heat diffusion equation 60
@@
As a representative example, Fig. 2(b)-(c) show and the (electron) temperature (11), respectively, calculated for Ag spheres of sizes up to nm dispersed in oil (permittivity , thermal conductivity Wm-1K-1) for an excitation laser wavelength close to the (dipolar) plasmon resonance ( nm) and for a wavelength farther from the resonance ( nm). The absorption cross-section (density) is well studied 61, hence, described below only briefly. Specifically, shows a peak around nm. For nm, the absorption cross-section (hence, ) takes the quasi-static form, (see Fig. 9 or more generally, Ref. [ 61]). As the sphere radius increases, the nm dipole resonance undergoes a red-shift (see Fig. 10(a)). Thus, since the chosen excitation wavelength nm is at the long wavelength tail of the resonance, increases with as the dipole resonance tunes towards . However, the strength of the dipole resonance decreases with growing sphere size, so that beyond nm, the absorption becomes limited to the surface as the importance of higher-order resonances increases. The corresponding size dependence of is shown in Fig. 9.
For excitation at nm, the averaged absorption cross-section density increases monotonically with the sphere size and its peak occurs beyond the range of consideration in this work, see Fig. 9(b). In that sense, the off-resonance illumination case behaves as in the small size regime of the on-resonance case; we will not dwell on it further in this study.

### v2

@@ 1 Introduction
The emission of light from metal nanostructures induced by illumination, frequently referred to colloquially as metal photoluminescence (PL), is a fundamental process occurring in a prototypical solid-state system111In many cases, the term photoluminescence is used to refer to spontaneous emission / radiative recombination of an electron and a hole, in contrast to inelastic light scattering, aka electronic Raman scattering. In that sense, in this work we use the term photoluminescence (PL) as a shorthand for emission, without distinguishing between these possibilities.. It also has practical importance - metal nanoparticles are frequently used as non-bleaching fluorescent labels in bio-imaging 1, 2, 3, correlative light-electron microscopy 4, and more recently, as enablers of thermometry 5, 6, 7, 8, 9.
@@
Despite that dual importance, it is surprising that many aspects of the metal PL have been under debate for decades. These debates are on, for example, whether the emission occurs due to recombination of an electron and hole residing both within the conduction band 10, 11, 12 or involving a hole in the valence band (e.g., Ref. [ 13, 14, 15, 16, 17, 18]), on whether the emission statistics is fermionic or bosonic (e.g., Ref. [ 7, 19]), thermal or non-thermal (e.g., Ref. [ 20, 21, 11, 22, 12]), or how it depends on the local electric field (e.g., Ref. [ 20, 21, 11, 23]). Of particular interest is the dependence on the size of the nanostructure (e.g., Ref. [ 24, 25, 26, 16, 17]). Many early works studied the emission from metal clusters or few nm particles, primarily focussing on atomistic considerations and ligand coverage; these works highlighted the trade-off between the presence of a plasmonic resonance and the quantum yield for emission, see Ref. [ 27]. For larger structures, there have been reports of contradicting trends. For example, Dulkeith et al. Ref. [ 24] showed that the (time and) frequency integrated Stokes emission (SE) from spheres of growing radii illuminated by a short pulse scales linearly with their volume; similar results were reported by Gaiduk et al. Ref. [ 25] for continuous wave (CW) illumination. On the other hand, Lin et al. Ref. [ 28] and more recently, Bowman et al. Ref. [ 29] have demonstrated that the CW PL from rods and thin Au films, respectively, decreases with growing thickness; similar trends were reported in Ref. [ 17, 30] for pulsed illumination. Other works reported a mixed trend (e.g., Ref. [ 26]).
@@
These debates originate from the different conditions under which the PL measurements have been made (in terms of the number, density and shape of the nanostructure geometry, the illumination duration, intensity and wavelength, the relative position of the latter with respect to the resonances of the structure, the relative spectral position of emission with respect to the illumination wavelength, the geometry and thermal properties of the surrounding etc.) or the exact quantity studied (temporally and/or spectrally resolved/integrated PL), but also from the absence of a simple theory that encompasses both photonic aspects of the problem (via the photonic density of states, as appearing in Purcell’s formula, see e.g., 22), as well as the thermal aspects and the electronic aspects, i.e., the distribution and associated weights of the continuum of possible recombination transitions.
@@
Recently, Dubi and Sivan employed a simple analytic expression for the steady-state electron distribution in a Drude metal to derive an equally simple expression for the PL from metals illuminated by CW light 12. By relying on the empiric values of the permittivity to determine the absorption, this approach circumvents the need to specify the transition matrix element and circumvents the arguments about the origins of the emission (being radiative recombination or electronic Raman / inelastic light scattering etc. 5, 31, 29, 32, 33, 34). This predicted behaviour was verified experimentally in Ref. [ 18], and enabled the resolution of many of the disagreements described above (in particular, associated with the emission statistics and electric-field dependence). This expression was more recently extended to explain the dependence of the PL following short pulse illumination on the electric field 35.
@@
In this work, we use the approach of Ref. [ 12] to reconcile the various seemingly contradicting reports on the size-dependence of the PL from the prototypical geometries of metal particles and films. Specifically, in Section 2, we combine the analytic expression of Ref. [ 12] with the detailed analysis of the heating of illuminated metal spheres 36 and most importantly, with recent progress made by Loirette-Pelous and Greffet who showed how to determine the total PL of a metal nanostructure in the case of a non-thermal distribution 37.
@@
In Section 3, we use the resulting analytic expression to compute the PL from metal nano-spheres and nano-films as a function of their size. We distinguish between three cases. For weak illumination (hence, negligible heating), the dependence of the PL spectrum (both the Stokes Emission (SE) and the anti-Stokes Emission (aSE)) on the structure size is determined by a single parameter - the absorption cross-section density (or equivalently, via Kirchhoff’s Law, the emission cross-section density); the electronic contribution is size-independent in this regime. Thus, generically, the PL scales with the illuminated volume, i.e., it is a volume effect for small sizes and it becomes a surface effect when the size exceeds the penetration depth of a few 10s of nm. However, in the latter case, the resonances characteristic of particles modify further the size-dependence of the PL and effectively dominate it. Indeed, a change of size causes the emission at a given frequency to shift in and then out of resonance.
@@
For stronger illumination (hence, moderate heating), the size-dependence of the PL is amplified through the dependence of the Bose function on the (electron) temperature. This has a fairly small effect on the SE, but a large effect on the aSE due to the exponential dependence of the Bose function on the temperature, which itself scales with the absorption cross-section density. For even stronger heating, thermo-optic effects kick in, and cause the quality factor of the resonance to decrease 38, 39. This causes weaker (excessive) heating at resonance (away from resonance), and hence, has a complex effect on the PL, depending on the emission frequency (SE/aSE).
@@
We then demonstrate a good qualitative match between the prediction of our model and experimental results for spheres and films; in fact, good quantitative matches are observed in most cases. We demonstrate different trends by looking also at other particle shapes (rods and nano disks), and find a good qualitative match, even in the presence of modest field and temperature non-uniformity levels. Remarkably, the agreement usually extends beyond the formal limits of the analysis, specifically, in the presence of interband emission events, which are not accounted for in our analytic expression.
@@ 2.1 A microscopic view
Based on the CW solution for (the steady-state electron distribution) obtained in Ref. [ 40, 41] for uniformly illuminated Drude metals and its experimental verification in Ref. [ 42, 43, 44, 45], in Ref. [ 12], Sivan and Dubi presented a quantitative theory for the PL from Drude metals under continuous wave (CW) illumination, showing that the probability of emitted photons per unit frequency is given approximately by
@@
represents the emission rate of a single electron. In the above, and are the emitter position vector and frequency, is the excitation frequency, is the Fermi energy, is the electron density of states and is the local field. is the transition dipole moment between electronic states with an initial energy and final energy (assumed to be energy-independent) and is the local density of photonic states (LDOPS).
The electronic contribution of the emission formula (Eq. (2)) was shown in Ref. [ 12] to consist of a series of Planck’s black-body radiation-like terms , i.e.,
@@
Here, is the (effective) electron temperature, extracted from the first energy moment of the electron distribution 46, 47, 40, 41; for CW illumination, the latter is nearly equal to the phonon temperature, however, since the PL literature is inconsistent on the matter, the notation adopted in this manuscript emphasizes that the dependence is on the electron temperature. Further, the absorbed power density () is defined in terms of the local as
@@
where is the electron density and is an average rate of collisions between electrons. For simplicity, we set in Eq. (7) to a value typical for a non-thermal electron, i.e., we use the Fermi Liquid theory expression 48, 40 at .
The first term on the RHS of Eq. (2.1) represents the average energy of thermal emission per electromagnetic mode (i.e., for vacuum electric fields); the next terms represent the non-thermal emission caused by deviations of the electron distribution from thermal equilibrium due to one photon absorption (1PA), two photon absorption (2PA) etc.. As discussed in Ref. [ 12, 35], for CW illumination, these terms are typically small compared to the thermal emission at mid-IR frequencies, but dominate the emission close to the illumination frequencies and above them 12, 35. All these details are depicted in the schematic representation (Fig. 1). The complete step structure of the non-thermal contributions was observed experimentally for CW illumination for the first time in Ref. [ 18].
Notably, the emission (1) is assumed to occur only within the conduction band, i.e., to involve only intraband transitions; in that sense, strictly speaking, should include only the intraband contribution (as in Ref. [ 37]). This would not matter for the illustrative examples below, for which we study Ag nanostructures, but could somewhat affect the match between our predictions and the experimental data (see Sections 3.1.4 and 3.2.3). In order to include interband absorption and emission transitions, our PL formula can be amended by accounting for the increased number of non-thermal holes above the interband absorption threshold in the electron distribution function , as shown in Fig. S10(b) of Ref. [ 29], and adding the valence bands to the electron density of states. As shown below, our analysis provides a qualitative (and usually, also a quantitative) match to the experimental data even without accounting for these additional transitions.
@@ 2.2 A macroscopic view
The result (1) applies for a general point within the metal. It was so far used to understand the general parametric dependence of the PL on the electric field, its statistics etc. with only a phenomenological address of the shape and size of the emitting nanostructure.
@@
In order to account for a specific nanostructure geometry, one needs to sum over the random spontaneous emission events from the nanostructure volume as a function of space. This should allow one to account also for the actual magnitude of the emission, thus, including also the portion of the emitted photons that got re-absorbed (sometimes referred to as re-cycled), rather than only the probability of emission (as given by (1)). This can be done using the local Kirchhoff Law for non-isothermal bodies 49, 50, 51 and its extension to metals having non-equilibrium electron distributions 37. The latter study showed that the actual emitted power in the direction into a solid angle (; see Eq. (9.15.9) in Ref. [ 52]) is given by
where is the photonic density of states of free space, and are the polarization of the emitted and incident waves, respectively, and is the emission cross-section density (aka emissivity density), which by the local Kirchhoff Law, equals the absorption cross-section density. In that regard, following Ref. [ 49, 50, 51], and unlike in Ref. [ 37], where the PL was expressed in terms of absorption cross-section, Eq. (8) utilizes the absorption cross-section density to extend the formulation to structures with nonuniform field distributions.
@@
Eq. (9) shows that the quantities that determine the emission are the absorption cross-section density and the (electron) temperature distribution. This dependence is similar to that provided in Ref. [ 53], as well as to the scaling of surface-enhanced Raman Scattering (see Ref. [ 54]), but goes beyond these theories by providing a rigorous (rather than an empiric 53) account of the effect of the electron distribution to the PL. As shown below, the expression (9) also circumvents the more advanced yet complicated -space calculations (as e.g., in Ref. [ 30, 29, 33]), without compromising much predictive capabilities.
@@
We now note that the spatially-integrated emission (9) involves space-varying functions. Specifically, the spatial distribution of is determined by the material constituents and geometry of the nanostructure. In addition to gradients induced by illumination non-uniformity (e.g., when a focussed beam illuminates a film), Eq. ((9)) exhibits significant gradients on scales exceeding the penetration (skin) depth, i.e., for more than a few tens of nm for noble metals. In contrast, the non-uniformity of temperature is typically much weaker than that of the electric field (and hence ) due to the strong (electron, hence) heat diffusion in metals. This (electron) temperature non-uniformity should, in principle, be extracted from a self-consistent solution of microscopic equations for the electron dynamics; however, in practice, since the deviation from thermal equilibrium is minute for CW illumination 40, 41, 45, it is usually determined by coarse-graining such equations into heat equations, see, e.g., Ref. [ 55, 47, 40, 41]. The resulting single or two temperature models reveal that heat gradients and electron-phonon temperature differences are very small (see, e.g., Ref. [ 56, 57, 58, 59, 60, 61]). Finally, like the absorption cross-section density and unlike the electron temperature , the non-thermal part of the electron distribution, manifested via (5), is roughly determined by the local electric field distribution. As predicted in Ref. [ 31, 62, 63] and demonstrated experimentally in Ref. [ 29], the reason for that is the minimal (few nm) mean free path of these electrons, which in turn, originates from their femtosecond-scale collision rate. Accordingly, to determine the spatial dependence of the PL integrand, in what follows we use heat equations to determine the electron temperature and rely on the solution of Maxwell’s equations to determine the electric field, and hence, and . Having said that, in many cases (specifically, for few nm spheres and uniformly illuminated thin films, both studied below), is quite uniform. In such cases, can be estimated from measurable quantities like the absorption cross-section, (in the case of particles) or absorptance, (in the case of films), thus, simplifying the formulation slightly, and making the analysis of the PL simpler.
@@ 3.1.1 Determination of the electron temperature
We start our analysis by considering the PL from nano-spheres under CW plane wave illumination in a uniform optical and thermal environment (oil, in our case), see Fig. 2(a). We choose silver as a prototypical plasmonic material, as it does not require accounting for interband transitions for illumination with visible light. As mentioned above, in this case, it is customary to neglect the small difference between the electron and phonon temperatures (see, e.g., Ref. [ 64, 40, 41]) and consider a single temperature model. Nevertheless, as mentioned, we choose to denote the temperature below as , in order to emphasize that it is the electron temperature that is the relevant quantity as far as photon emission in the visible and near infrared spectral regimes is concerned. Then, the steady-state temperature can be obtained by solving the heat diffusion equation 65
@@
As a representative example, Fig. 2(b)-(c) show and the (electron) temperature (11), respectively, calculated for Ag spheres of sizes up to nm dispersed in oil (permittivity , thermal conductivity Wm-1K-1) for an excitation laser wavelength close to the (dipolar) plasmon resonance ( nm) and for a wavelength farther from the resonance ( nm). The absorption cross-section (density) is well studied 66, hence, described below only briefly. Specifically, shows a peak around nm. For nm, the absorption cross-section (hence, ) takes the quasi-static form, (see Fig. LABEL:Fig.ACS_vs_radius or more generally, Ref. [ 66]). As the sphere radius increases, the nm dipole resonance undergoes a red-shift (see Fig. LABEL:Fig.alpha_vs_WL(a)). Thus, since the chosen excitation wavelength nm is at the long wavelength tail of the resonance, increases with as the dipole resonance tunes towards . However, the strength of the dipole resonance decreases with growing sphere size, so that beyond nm, the importance of higher-order resonances increases, and thus, the absorption becomes limited to the surface. The corresponding size dependence of is shown in Fig. LABEL:Fig.ACS_vs_radius.
Similarly, for excitation at nm, the averaged absorption cross-section density increases monotonically with the sphere size for all sizes studied here since the peak response occurs beyond the range of consideration in this work, see Fig. LABEL:Fig.ACS_vs_radius(b). In that sense, the off-resonance illumination case behaves as in the small size regime of the on-resonance case; we will not dwell on it further in this study.